# Quality-Aware Translation Models: Efficient Generation and Quality Estimation in a Single Model

### v3

@@ 2 Related Work
While conventional MT research often relies on MAP decoding or generating k-best lists through beam search for MBR decoding, Eikema and Aziz (2020) proposed an approximation of MBR decoding via unbiased sampling. Their method aims to address the limitations of MAP decoding (Eikema and Aziz, 2020; Müller and Sennrich, 2021; Eikema and Aziz, 2022) by demonstrating that samples drawn from the NMT model align more faithfully with training data statistics when compared to beam search. Freitag et al. (2022a) showed that using neural metrics instead of overlap metrics results in significant improvements in translation quality. As a follow up, Freitag et al. (2023a) reported that the choice of sampling approach is important and epsilon sampling (Hewitt et al., 2022) is ideal for MBR decoding and reranking. Cheng and Vlachos (2023) introduced an orthogonal method to our proposed approach. They speed up MBR decoding by increasing the number of samples gradually used to estimate the utility while pruning hypotheses.
@@ 3 Preliminaries
In many applications we want to generate diverse hypothesis, e.g. in generative tasks where creativity is desired. In this case, instead of selecting the candidate with the highest probability (or an approximation thereof), we sample the output sentence following the probability distribution defined by the model. For NMT, this approach is used for generating a list of candidate translations, e.g. for MBR decoding. Specifically, epsilon sampling, as outlined by Hewitt et al. (2022), has emerged as the leading sampling technique for MBR. It was shown by Freitag et al. (2023a) to outperform other methods such as ancestral, top-k or nucleus sampling (Holtzman et al., 2020). Epsilon sampling prunes away any token with a probability lower than a given threshold $\varepsilon$, thereby guaranteeing that each token within a sample is allocated a fair probability mass. The likelihood of selecting token $y^{(\tau)}$ in the sampling process at time $\tau$ is governed by
@@ 3.2 Minimum Bayes Risk Decoding
$h^{\rm best}$ $=$ $\operatorname*{arg\,max}_{y\in{\cal Y}}\sum_{r\in\Omega}u(y,r)P_{\rm human}(r|x),$ (2)
@@
$h^{\rm MBR}=\operatorname*{arg\,max}_{y\in{\cal Y}}\frac{1}{|{\cal Y_{\rm model}}|}\sum_{r\in\cal Y_{\rm model}}u(y,r),$ (3)
@@ 4.1.1 Assessing Quality
We first prepare the training dataset by computing the translation quality of each training sample and labelling each sentence pair with the corresponding quality score. Since the distribution over translation qualities is not necessarily uniform, we discretize the scores via equal mass binning into $B$ bins, which are then mapped to single tokens of the vocabulary of the translation model. This binning strategy leads to a balanced training set w.r.t. quality scores. To achieve this, we consider the set of quality scores $Q$ from all samples in the training dataset, denoted as ${q_{1},q_{2},\ldots,q_{N}}$, in order to determine bin boundaries or cut-off points ${c_{1},c_{2},\ldots,c_{B+1}}$ such that
|$\left|\{x\;|\;q\in Q,\;c_{i}\leq q<c_{i+1}\}\right|\approx\frac{N}{B}\qquad\forall 1\leq i\leq B.$ (4)
In this way each bin contains approximately the same number of samples. This is to avoid sample imbalances per bin when training the model, which in preliminary experiments proved to be important as to not bias the model towards the most frequent label. Next, each bin is assigned a bin identifier $b$, which is represented by a single token in the model vocabulary. E.g. if we define 10 bins22210 bins is the number we use in our experiments. For an exploration on other number of bins see Appendix D, we can just use the numbers between 0 and 9. Lastly, the quality score is inserted into the data pipeline during training to associate the source and target pair with the respective bin identifier $b$. To mark the token $b$ as a QE value, we employ a special string format by surrounding $b$ with square brackets: [$b$]. In the following we outline two distinct methods for integrating our quality score string into the model.
@@ 4.1.3 Quality-Aware Prediction (QA Prediction)
If using a reranking approach these quality scores can be used to reorder samples within a candidate list. However, due to our use of discretized bins, it is possible that the model predicts the same scores for multiple samples. To address this, we also consider the log probabilities $z$ associated with the bin identifier tokens. This additional information allows for a more precise reranking of samples in the candidate list. Specifically, we sort samples with respect to $b$ as the primary sorting criterion and use the log probabilities $z$ as the secondary criterion. Given a candidate list of size $M$, we sort the samples into ${y_{1},y_{2},\ldots,y_{M}}$ with corresponding discretized quality scores ${b_{1},b_{2},\ldots,b_{M}}$ and log probabilities ${z_{1},z_{2},\ldots,z_{M}}$ in such manner that
$\end{multlined}\forall 1\leq i<j\leq M:(b_{i}>b_{j})\;\text{or}\;(b_{i}=b_{j}\;\text{and}\;z_{i}>z_{j})$ (5)
@@ 5 Experimental Setup
Our model is based on a transformer architecture, consisting of 6 encoder and 6 decoder layers, 16 attention heads with a dimension of 128, a hidden dimension of 8192, and a model dimension of 1024, resulting in 551M parameters. In our experiments, we employ a shared vocabulary of 32k tokens and impose a maximum sentence length of 128 tokens. We utilize GELUs with gated activation functions. The baseline system is trained on the entire available dataset. All models are trained on TPUs (v3) until they reach convergence. The MetricX-QE model used for quality estimation is a transformer based model with a total of 2B parameters, as described in Juraska et al. (2023). To assess the applicability of our approach to LLMs, we also fine-tune and evaluate a quality-aware LLM, with results in Appendix C.
@@
We use neural metrics for evaluation, with a focus on Comet (Rei et al., 2020a) (Comet 22 version). We also report MetricX scores, but as MetricX-QE is based on it and our methods directly optimize this metric, there is the danger of overfitting for this particular metric (Amrhein and Sennrich, 2022; Yan et al., 2023). In addition, for selected experiments we conducted additional expert-based human evaluations using MQM (Freitag et al., 2021), a human evaluation scheme centered on marking errors present in the translations. Details can be found in Appendix I. For completeness we also report Bleu scores in Appendix B, but we do not analyze them here, as it has been shown that they do not correlate well with human judgement of NMT systems Freitag et al. (2022b).
@@ 6.1.1 Can NMT models learn to estimate quality?
In a first experiment, we train a model using the proposed Quality-Aware Prediction method (§ 4.1.3). We then use this model to force-decode the translations provided in the WMT23 QE shared task Blain et al. (2023) and, once the model has reached the end of the translation, we select the quality label that gets assigned the highest probability by the model. As many different segments will be assigned the same discrete quality label, we also use model probability to break ties when calculating correlations (see Section 4.1.3). In this way we can check if our translation model can double as a QE model. We evaluate this approach using the same metrics as in the shared task. The results are shown in Table 1, together with some other representative metrics. The top block in Table 1 shows the top-scoring metrics in the shared task. [^3] Note that evalation metrics and QE metrics were evaluated simultaneously. As expected, we see that our model does not achieve the performance of separate dedicated models (and specifically not that of the MetricX-QE metric it is based on), but it still outperforms the traditional Bleu and ChrF metrics, and these have access to the reference translations.
@@ 6.1.2 Can NMT models distinguish the quality of their own outputs?
Next we turn to the question of whether the system is able to judge the quality of its own translations. In order to test this, we translated our dev dataset with quality label predictions and then computed the actual MetricX-QE scores which the model is trained to predict. We find that the predicted quality score bins are well aligned with the ground truth scores, as demonstrated in Figure 1, where we show the distribution of ground truth scores (non discretized) across the predicted bins. In particular we note that the system is able to accurately detect bad quality translations, which will become important for the translation use case.
Figure 1: Alignment between predicted quality scores from the QA Prediction model and ground truth MetricX-QE scores of translations in the en $\rightarrow$ de test dataset. The boxplots show the distribution of ground truth scores across all samples assigned to each bin. The median ground truth quality score increases steadily in line with the predicted bins.
@@ 6.1.3 Can we control translation quality?
As a last control experiment we train a QA-prompting model following Section 4.1.2 and generate translations using different quality labels. I.e. we are asking the model to generate good translation (high scores in the quality labels) as well as bad translations (low scores). The results are shown in Figure 2. It can be seen that the system is indeed able to adjust the quality of the translation output according to the given prompting. Example translation outputs are shown in Appendix F.
@@ 6.2 Translation Performance
Baseline Full 80.2 85.8 1.81
QA Prompting (Ours) Full 82.3 87.1  ${}^{\star}$ 1.43  ${}^{\star}$
QA Prediction (Ours) Full 82.0 86.5  ${}^{\star}$ 2.07
External QE-Reranking Full 83.3 86.9  ${}^{\star}$ 1.50  ${}^{\star}$
Baseline Filter 81.8 87.0
QA Prompting (Ours) Filter 82.6 87.3  ${}^{\star}$
QA Prediction (Ours) Filter 82.5 86.7
External QE-Reranking Filter 83.7 86.9
Table 2: Comparison between quality-aware models and baseline models on the full and filtered training datasets. The quality-aware methods outperform the baseline model and perform similar to reranking without requiring an additional MetricX-QE model during decoding. ${}^{\star}$ denotes statistically significant (pairwise permutation test Koehn (2004) with p=0.05) differences compared to the baseline with $p<0.05$. No significance is computed for MetricX due to the methods optimizing this metric directly.
@@
(a) MetricX- Filtered data (b) Comet- Filtered data
@@
Table 2 shows that both quality-aware methods surpass the baseline model in terms of MetricX and Comet, with QA Prompting showing better results on both metrics. As expected, all the methods achieve big improvements in MetricX, as it is closely related to the MetricX-QE metric that we are directly optimizing. In fact, external QE-Reranking achieves the best MetricX score by a wide margin, however Comet puts QA Prompting on-par with QE-Reranking.
@@
When compared to the stronger baseline with training data prefiltered with MetricX-QE scores, we see that QA-prompting is still able to obtain a slight improvement over both the baseline and QE-Reranking, which is still statistically significant. QA-prediction is not effective in this setting either.
@@
External QE-Reranking 83.3 86.9  ${}^{\star}$ 1.50  ${}^{\star}$
$M=50$ MBR Baseline 82.5 86.8  ${}^{\star}$ 1.41  ${}^{\star}$
MBR QA Prompting (Ours) 83.5 87.4   ${}^{\star{\dagger}}$ 1.36   ${}^{\star{\dagger}}$
MBR QA Prediction (Ours) 83.4 87.5   ${}^{\star{\dagger}}$ 1.45   ${}^{\star{\dagger}}$
@@
MBR QA Prompting (Ours) 82.6 86.9  ${}^{\star}$
MBR QA Prediction (Ours) 82.8 87.0  ${}^{\star}$
@@
MBR QA Prompting (Ours) 83.8 87.7   ${}^{\star{\dagger}}$
MBR QA Prediction (Ours) 83.9 87.7   ${}^{\star{\dagger}}$
@@
MBR QA Prompting (Ours) 83.1 87.3  ${}^{\star}$
MBR QA Prediction (Ours) 83.1 87.2  ${}^{\star}$
Table 3: MBR results with Quality-Aware decoding approached. ${}^{\star}$ denotes statistically significant differences compared to the baseline with $p<0.05$, ${}^{\dagger}$ denotes statistically significant differences compared to the MBR baseline with $M=50$. No significance is computed for MetricX due to the methods optimizing this metric directly.
@@ 6.3 MBR decoding
For baseline MBR decoding we use epsilon sampling to generate a candidate list of $M$ samples, similarly for the Quality Aware prompting approach. For Quality-Aware Prediction, in line with the previous section, we employ our Quality-Aware Prediction approach to sample $N=1024$ hypotheses with quality score strings. Then we rank all samples and select the top $M$ samples as our candidate list. Subsequently, MBR decoding is applied to the gathered candidate lists. We used Bleurt as utility function as a proxy for MetricX, due to the high computational cost of this last metric.
In Figure 3 we show the performance of our quality-aware approaches compared to baseline MBR decoding across various candidate list sizes $M$. Our proposed methods consistently outperform baseline MBR decoding in terms of MetricX and Comet scores, irrespective of the candidate list size. Notably, our quality-aware approaches combined with MBR decoding require substantially fewer candidates to achieve equivalent performance to baseline MBR decoding. For example, in models trained on the entire dataset, the Quality-Aware Prompting and Quality-Aware Prediction approaches obtain Comet scores of 86.9 and 87.0, respectively, with a candidate list size of 5 (with just 20 utility function computations [^4] The number of computations is $M\times(M-1)$ as a hypothesis is not evaluated against itself. per sentence). In contrast, baseline MBR decoding plateaus at 87 Comet starting at a candidate size of 50 (requiring 2450 utility function computations per sentence). This translates to a 100-fold increase in computations for the baseline model to achieve a similar score. We also note that the baseline model with a candidate list size of 1024 achieves a MetricX score of 82.9 and a Comet score of 87.0. This indicates that our approach, with a candidate list size of 50, outperforms a baseline model with even 1024 samples. For our experiments on filtered data we observe a similar improvement in the quality-aware models when compared to baseline MBR decoding.

### v4

@@ 2 Related Work
While conventional MT research often relies on MAP decoding or generating k-best lists through beam search for MBR decoding, Eikema and Aziz (2020) proposed an approximation of MBR decoding via unbiased sampling. Their method aims to address the limitations of MAP decoding (Eikema and Aziz, 2020; Müller and Sennrich, 2021; Eikema and Aziz, 2022) by demonstrating that samples drawn from the NMT model align more faithfully with training data statistics when compared to beam search. Freitag et al. (2022a) showed that using neural metrics instead of overlap metrics results in significant improvements in translation quality. As a follow up, Freitag et al. (2023a) reported that the choice of sampling approach is important and epsilon sampling (Hewitt et al., 2022) is ideal for MBR decoding and reranking. Cheng and Vlachos (2023) introduced an orthogonal method to our proposed approach. They speed up MBR decoding by gradually increasing the number of samples used to estimate the utility while pruning hypotheses.
@@ 3 Preliminaries
In many applications we want to generate diverse hypotheses, e.g. in generative tasks where creativity is desired. In this case, instead of selecting the candidate with the highest probability (or an approximation thereof), we sample the output sentence following the probability distribution defined by the model. For NMT, this approach is used for generating a list of candidate translations, e.g. for MBR decoding. Specifically, epsilon sampling, as outlined by Hewitt et al. (2022), has emerged as the leading sampling technique for MBR. It was shown by Freitag et al. (2023a) to outperform other methods such as ancestral, top-k or nucleus sampling (Holtzman et al., 2020). Epsilon sampling prunes away any token with a probability lower than a given threshold $\varepsilon$, thereby guaranteeing that each token within a sample is allocated a fair probability mass. The likelihood of selecting token $y^{(\tau)}$ in the sampling process at time $\tau$ is governed by
@@ 3.2 Minimum Bayes Risk Decoding
$h^{\rm best}=\operatornamewithlimits{argmax}_{y\in{\cal Y}}\sum_{r\in\Omega}u(y,r)P_{\rm human}(r|x),$ (2)
@@
$h^{\rm MBR}=\operatornamewithlimits{argmax}_{y\in{\cal Y}}\frac{1}{|{\cal Y_{\rm model}}|}\sum_{r\in\cal Y_{\rm model}}u(y,r),$ (3)
@@ 4.1.1 Assessing Quality
We first prepare the training dataset by computing the translation quality of each training sample and labelling each sentence pair with the corresponding quality score. Since the distribution over translation qualities is not necessarily uniform, we discretize the scores via equal mass binning into $B$ bins, which are then mapped to single tokens of the vocabulary of the translation model. This binning strategy leads to a balanced training set w.r.t. quality scores. To achieve this, we consider the set of quality scores $Q$ from all samples in the training dataset, denoted as ${q_{1},q_{2},\ldots,q_{N}}$, in order to determine bin boundaries or cut-off points ${c_{1},c_{2},\ldots,c_{B+1}}$ such that $\forall 1\leq i\leq B$
|$\left|\{x\;|\;q\in Q,\;c_{i}\leq q<c_{i+1}\}\right|\approx\frac{N}{B}.$ (4)
In this way each bin contains approximately the same number of samples. This is to avoid sample imbalances per bin when training the model, which in preliminary experiments proved to be important as to not bias the model towards the most frequent label. Next, each bin is assigned a bin identifier $b$, which is represented by a single token in the model vocabulary. E.g. if we define 10 bins (the actual number we used in our experiments [^2] For an exploration on other number of bins see Appendix D.), we can just use the numbers between 0 and 9. Lastly, the quality score is inserted into the data pipeline during training to associate the source and target pair with the respective bin identifier $b$. To mark the token $b$ as a QE value, we employ a special string format by surrounding $b$ with square brackets: [$b$]. In the following we outline two distinct methods for integrating our quality score string into the model.
@@ 4.1.3 Quality-Aware Prediction (QA Prediction)
If using a reranking approach these quality scores can be used to reorder samples within a candidate list. However, due to our use of discretized bins, it is possible that the model predicts the same scores for multiple samples. To address this, we also consider the log probabilities $z$ associated with the bin identifier tokens. This additional information allows for a more precise reranking of samples in the candidate list. Specifically, we sort samples with respect to $b$ as the primary sorting criterion and use the log probabilities $z$ as the secondary criterion. Given a candidate list of size $M$, we sort the samples into ${y_{1},y_{2},\ldots,y_{M}}$ with corresponding discretized quality scores ${b_{1},b_{2},\ldots,b_{M}}$ and log probabilities ${z_{1},z_{2},\ldots,z_{M}}$ such that
[…]
@@ 5 Experimental Setup
Our model is a transformer consisting of 6 encoder and 6 decoder layers, 16 attention heads with a dimension of 128, a hidden dimension of 8192, and a model dimension of 1024, resulting in 551M parameters. We employ a shared vocabulary of 32k tokens and impose a maximum sentence length of 128 tokens. We utilize GELUs with gated activation functions. The baseline system is trained on the entire available dataset. All models are trained on TPUs (v3) until they reach convergence. The MetricX-QE model used for quality estimation is a transformer based model with a total of 2B parameters, as described in Juraska et al. (2023). To assess the applicability of our approach to LLMs, we also fine-tune and evaluate a quality-aware LLM, with results in Appendix C.
@@
We use neural metrics for evaluation, with a focus on Comet (Rei et al., 2020a) (Comet 22 version). We also report MetricX scores, but as MetricX-QE is based on it and our methods directly optimize this metric, there is the danger of overfitting for this particular metric (Amrhein and Sennrich, 2022; Yan et al., 2023). In addition, for selected experiments we conduct expert-based human evaluations using MQM (Freitag et al., 2021), a human evaluation scheme centered on marking errors present in the translations. Details can be found in Appendix I. For completeness we also report Bleu scores in Appendix B, but we do not analyze them here, as it has been shown that they do not correlate well with human judgement of NMT systems Freitag et al. (2022b).
@@ 6.1.1 Can NMT Models Learn to Estimate Quality?
In a first experiment, we train a model using the proposed Quality-Aware Prediction method (§ 4.1.3). We then use this model to force-decode the translations provided in the WMT23 QE shared task Blain et al. (2023) and, once the model has reached the end of the translation, we select the quality label that gets assigned the highest probability by the model. As many different segments will be assigned the same discrete quality label, we also use model probability to break ties when calculating correlations (see Section 4.1.3). In this way we can check if our translation model can double as a QE model. We evaluate this approach using the same metrics as in the shared task. The results are shown in Table 1, together with some other representative metrics. The top block in Table 1 shows the top-scoring metrics in the shared task. [^3] Note that evaluation metrics and QE metrics were evaluated simultaneously. As expected, we see that our model does not achieve the performance of separate dedicated models (and specifically not that of the MetricX-QE metric it is based on), but it still outperforms the traditional Bleu and ChrF metrics, and these have access to the reference translations.
@@ 6.1.2 Can NMT Models Distinguish the Quality of Their Own Outputs?
Next we turn to the question of whether the system is able to judge the quality of its own translations. In order to test this, we translated our dev dataset with quality label predictions and then computed the actual MetricX-QE scores which the model is trained to predict. We find that the predicted quality score bins are well aligned with these “ground truth” scores444“Ground truth” in this context as these are the scores that the model was trained to predict. They are not necessarily ground truth for true quality measurement., as demonstrated in Figure 1, where we show the distribution of ground truth scores (non discretized) across the predicted bins. In particular we note that the system is able to accurately detect bad quality translations, which will become important for the translation use case.
Figure 1: Alignment between predicted quality scores from the QA Prediction model and actual MetricX-QE scores of translations in the en $\rightarrow$ de test dataset. The boxplots show the distribution of actual scores across all samples assigned to each bin. The median ground truth quality score increases steadily in line with the predicted bins.
@@ 6.1.3 Can We Control Translation Quality?
As a last control experiment we train a QA Prompting model following Section 4.1.2 and generate translations using different quality labels. I.e. we are asking the model to generate good translations (high scores in the quality labels) as well as bad translations (low scores). The results are shown in Figure 2. It can be seen that the system is indeed able to adjust the quality of the translation output according to the given prompting. Example translation outputs are shown in Appendix F.
@@ 6.2 Translation Performance
Baseline Full Corpus 80.2 85.8 1.81
QA Prompting (Ours) Full Corpus 82.3 87.1   1.43
QA Prediction (Ours) Full Corpus 82.0 86.5   2.07
External QE-Reranking Full Corpus 83.3 86.9   1.50
Baseline Filtered 81.8 87.0
QA Prompting (Ours) Filtered 82.6 87.3
QA Prediction (Ours) Filtered 82.5 86.7
External QE-Reranking Filtered 83.7 86.9
Table 2: Comparison between quality-aware models and baseline models on the full and filtered training datasets. The quality-aware methods outperform the baseline model and perform similar to reranking without requiring an additional MetricX-QE model during decoding. denotes statistically significant (pairwise permutation test Koehn (2004) with p=0.05) differences compared to the baseline with $p<0.05$. No significance is computed for MetricX due to the methods optimizing this metric directly.
@@
(c) MetricX- Filtered data (d) Comet- Filtered data
@@
Table 2 shows that both quality-aware methods surpass the baseline model in terms of MetricX and Comet, with QA Prompting showing better results on both metrics. As expected, all the methods achieve big improvements in MetricX, as it is closely related to the MetricX-QE metric that we are directly optimizing. In fact, external QE-Reranking achieves the best MetricX score by a wide margin, however Comet puts QA Prompting on-par with QE-Reranking. Even though an improvement of COMET of 0.2-0.4 seems small, Lo et al. (2023) use human judgment to verify differences in COMET scores and conclude that even small differences in COMET scores can mean large quality improvements.
@@
When compared to the stronger baseline with training data prefiltered with MetricX-QE scores, we see that QA Prompting is still able to obtain a slight improvement over both the baseline and QE-Reranking, which is still statistically significant. QA Prediction is not effective in this setting either.
@@
External QE-Reranking 83.3 86.9   1.50
$M=50$ MBR Baseline 82.5 86.8   1.41
MBR QA Prompting (Ours) 83.5 87.4   ⋆† 1.36   ⋆†
MBR QA Prediction (Ours) 83.4 87.5   ⋆† 1.45   ⋆†
@@
MBR QA Prompting (Ours) 82.6 86.9
MBR QA Prediction (Ours) 82.8 87.0
@@
MBR QA Prompting (Ours) 83.8 87.7   ⋆†
MBR QA Prediction (Ours) 83.9 87.7   ⋆†
@@
MBR QA Prompting (Ours) 83.1 87.3
MBR QA Prediction (Ours) 83.1 87.2
Table 3: MBR results with quality-aware decoding approaches. The symbol denotes statistically significant differences compared to the baseline with $p<0.05$, denotes statistically significant differences compared to the MBR baseline with $M=50$. No significance is computed for MetricX due to the methods optimizing this metric directly.
@@ 6.3 MBR decoding
For baseline MBR decoding we use epsilon sampling to generate a candidate list of $M$ samples, as we do for the Quality-Aware Prompting approach. For Quality-Aware Prediction, in line with the previous section, we employ our Quality-Aware Prediction approach to sample $N=1024$ hypotheses with quality score strings. Then we rank all samples and select the top $M$ samples as our candidate list. Subsequently, we apply MBR decoding to the gathered candidate lists. We used Bleurt as utility function as a proxy for MetricX, due to the high computational cost of this last metric.
In Figure 3 we show the performance of our quality-aware approaches compared to baseline MBR decoding across various candidate list sizes $M$. Our proposed methods consistently outperform baseline MBR decoding in terms of MetricX and Comet scores, irrespective of the candidate list size. Notably, our quality-aware approaches combined with MBR decoding require substantially fewer candidates to achieve equivalent performance to baseline MBR decoding. For example, in models trained on the entire dataset, the Quality-Aware Prompting and Quality-Aware Prediction approaches obtain Comet scores of 86.9 and 87.0, respectively, with a candidate list size of 5 (needing just 20 utility function computations [^5] The number of computations is $M\times(M-1)$ as a hypothesis is not evaluated against itself. per sentence). In contrast, baseline MBR decoding plateaus at 87 Comet starting at a candidate size of 50 (requiring 2450 utility function computations per sentence). This translates to a 100-fold increase in computations for the baseline model to achieve a similar score. We also note that the baseline model with a candidate list size of 1024 achieves a MetricX score of 82.9 and a Comet score of 87.0. This indicates that our approach, with a candidate list size of 50, outperforms a baseline model with even 1024 samples. For our experiments on filtered data we observe a similar improvement in the quality-aware models when compared to baseline MBR decoding.